OpenStudy (anonymous):

Factor as indicated: radical(x^2+1)-x^2/radical(x^2+1). One of the factors must be 1/radical(x^2+1).

OpenStudy (anonymous):

\[\sqrt{x^2+1}-\frac{x^2}{\sqrt{x^2-1}}\] like that?

OpenStudy (anonymous):

if you have to have one factor as \[\frac{1}{\sqrt{x^2+1}}\] then you have to factor it out \[\frac{1}{\sqrt{x^2+1}}\left(x^2+1-x^2\right)\]

OpenStudy (anonymous):

the point being that you had \[\frac{1}{\sqrt{x^2+1}}\] to begin with

OpenStudy (anonymous):

Yes, like that. That's what I got, but on my assignment it has that and then parentheses next to it, indicating another factor. So I guess I just put a one in the other parentheses because it's a trick question?